OpenStudy (anonymous):

What's the square root of 7/20? Or is there a site that can help me do fraction square roots?

OpenStudy (anonymous):

go to Google type in square root calculator and you should have it but the answer is approx 2.65

OpenStudy (phi):

you need a calculator if you want a decimal form of the answer. You can also do this \[ \sqrt{\frac{7}{20}}= \frac{\sqrt{7}}{\sqrt{20}} \] we can write 20 as \( 2\cdot 2 \cdot 5\) \[ \frac{\sqrt{7}}{\sqrt{20}} = \frac{\sqrt{7}}{\sqrt{ 2\cdot 2 \cdot 5}} \] notice we can "pull out" the pair of 2's from the bottom, and write your problem as \[ \frac{\sqrt{7}}{\sqrt{ 2\cdot 2 \cdot 5}}= \frac{\sqrt{7}}{2\sqrt{ 5}} \]

OpenStudy (phi):

which is not really much simpler. But it is a little bit simpler.

OpenStudy (anonymous):

Whoa, thanks!